OpenStudy (imkurstin.19):

Karen bakes cookies. She has observed that the number of cookies that turn out well in each batch varies with the size of the batch and is given by the function f(x) = 2x – 23, where x is the size of the batch (number of cookies in each batch). She never bakes fewer than 12 cookies in a batch. If the size of five batches is {20, 25, 26, 27, 34}, how many cookies turn out well in each batch?

OpenStudy (518nad):

okay so its telling you the size of each batch, we need so see how many turns out good for each one by plugging it into the equation

OpenStudy (518nad):

f(x) tells you the number of good ones for each x value

OpenStudy (518nad):

x= {20, 25, 26, 27, 34} I will do a couple of them for you f(x)=2x-23 f(20) = 2*20 -23 = 40-23 = 17 17 turn out good for a batch of 20 f(x) =2x-23 f(25)=2*25-23=50-23=27 so 27 turns out good for a batch of 27 (note**kind of silly because thats more than what she baked, the function is just bad) you can try for x=26,27,and 34